OpenStudy (anonymous):

Find the number c that satisfies the conclusion of the Mean Value Theorem. F(x)=x^3+x-8 [0,2]

OpenStudy (zarkon):

solve\[\frac{F(2)-F(0)}{2-0}=F'(c)\] for c

OpenStudy (anonymous):

-8-2/-2

OpenStudy (anonymous):

=5

OpenStudy (anonymous):

so now i find the derivative of the equation and put =5?

OpenStudy (zarkon):

5 is correct but you seemed to put it in backwards from what i wrote. yes...set the derivative of F equal to 5 and solve

OpenStudy (anonymous):

is the derivative is 3x^2+1 ?

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

i dont know what to do from there

OpenStudy (zarkon):

solve \(3x^2+1=5\)

OpenStudy (anonymous):

\[2/\sqrt{3}\] !!! thank you